SUB-DOPPLER SATURATION SPECTROSCOPY WITH THE COLOGNE THZ SPECTROMETER

\underline{S.P. BELOV}, G. WINNEWISSER and TH. KLAUS, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, D-50937 Cologne, Germany .

With the Cologne terahertz spectrometer (1) we have performed Lamp-dip measurements in the submillimeter wave region. The large output power of the frequency and phase stabilized backward wave oscillators (typically between 10 mW and 1 mW in the frequency range 300 GHz to 1000 GHz) was sufficient to saturate various molecular transitions. Together with the power reflected from the bolometer it has been possible to observe saturation dips in a free space absorption cell, whereby no further means have been required. Different molecules such as NH_{3}, HCl, and some of their isotopomers have been measured with sub-Doppler resolution at 570 GHz and 625 GHz. In the case of ammonia the ^{14}N hyperfine structure of the lowest ro-inversion transitions J = 1 ← 0 in the ground and J = 0 ← 1 in the v_{2} = 1 vibrational state has been fully resolved. The absolute frequencies have been measured with an accuracy of about 1 kHz. For the v_{2} = 1 state the quadrupole coupling constant eqQ and the nuclear spin-rotation coupling constant C_{N} determined from this saturation dip measurements are eqQ(v_{2}, J, K, s) = eqQ(1, 1, 0, s) = −4433.8(20) kHz and C_{N} = 6.38(60) kHz. The value of C_{N} for the vibrational ground state has been determined to be 6.7(3) kHz, whereby the eqQ value has been fixed to the number given by Hougen (2). This value is in very good agreement with the value C_{N} = 6.80734 kHz determined by Kukolich from two cavity maser measurements of the pure inversion transitions (3). The present data confirm new possibilities for high resolution molecular spectroscopy in the submillimeter wave region. The parameters of the spectrometer and some recent results of sub-Doppler saturation measurements will be discussed.

(1) G. Winnewisser, Vib. Spectrosc. \underline{8}, 241 (1995).

(2) J.T. Hougen, J. Chem. Phys. \underline{57}, 4207 (1972).

(3) S.G. Kukolich, Phys. Rev. \underline{156}, 83 (1967).